OpenStudy (anonymous):

Person will get medal and fan if they help. *NEED STEP BY STEP INSTRUCTIONS* Joe deposits $1,500 in an account that pays 3% annual interest compounded continuously. a. How much will Joe have in his account after 5 years? b. How long will it take Joe to double his money? Use natural logarithms and explain your answer.

OpenStudy (blurbendy):

Use the formula: A = Pe^(rt) Where P is the amount you start with (1,500) r is the rate (.05) in this case and t is time (after 5 years so t = 5) Can you solve question a from here?

OpenStudy (blurbendy):

Sorry the rate is (.03)

OpenStudy (anonymous):

yes hold on

OpenStudy (anonymous):

So this is the equation. 1500e^1.5?

OpenStudy (blurbendy):

e^.15 (.03 * 5) = .15

OpenStudy (anonymous):

For a. i got 1500e^1.5 = 6722.53

OpenStudy (blurbendy):

Nope, you should have gotten 1742.75 You know what e is on your calculator right?

OpenStudy (anonymous):

yeah.

OpenStudy (blurbendy):

I entered it like this into the calculator 1500e^(.03*5)

OpenStudy (anonymous):

Ok now i got it. so it is 1742.75

OpenStudy (blurbendy):

it's like 1742.751364 (I don't know how much you need to round).

OpenStudy (anonymous):

Its just to the tenth. now can you help me with part b? That is the harder part

OpenStudy (blurbendy):

Sure

OpenStudy (blurbendy):

This time we know what we need to get to double our money. 1500 * 2 = 3000 So put 3000 in for A in the equation we used before 3000 = 1500e(.03t) What we don't know is time. So we will solve for it. 3000 / 1500 = e^.03t 2 = e^.03t Now take the natural logarithm ON BOTH SIDES to get rid of e ln(2) = ln(e)^.03t This is important because ln(e) = 1 !!!! ln(2) = .03t ln(2) / 3 = t t = 23.1 years Understand?

OpenStudy (blurbendy):

The first part should read 3000 = 1500e^(.03t) I forgot to raise e to the power, but the rest is correct.

OpenStudy (anonymous):

THANK YOU SO MUCH!

OpenStudy (blurbendy):

No probs! Anytime :)